OpenStudy (impracticaljoker):

Find the ordered pairs for the x- and y-intercepts of the equation 5x − 6y = 30 and select the appropriate option below. (5 points) The x-intercept is (6, 0), the y-intercept is (0, −5). The x-intercept is (0, 6); the y-intercept is (−5, 0). The x-intercept is (5, 0); the y-intercept is (0, −6). The x-intercept is (0, 5); the y-intercept is (−6, 0).

OpenStudy (impracticaljoker):

@Will.H

OpenStudy (will.h):

set the equation in slope intercept form which means y isolated. -6y = -5x + 30 y = 5/6 x - 5 Y intercept is when you set x = 0 so y = (0,-5) and x intercept when you solve for x so x = (6,0)

OpenStudy (will.h):

Make sense?